lock:

Calculate the circle circumfrance's upper bound if the area is 400cm(square)

lock:

@Shadow

lock:

@theDeviliscoming

Shadow:

is there a picture

lock:

nope

lock:

The only given thing is 400 cm^2 area And they want you to calculate it in upper bound and lower bound..?

lock:

they didnt give me the raduis though

lock:

nvm i think its r≈11.28

lock:

so is it 11.20 and 11.30? @Shadow

jhonyy9:

do you know formula for circumference of a circle ?

jhonyy9:

area = 400 cm^2 you need to know formula of area of a circle too

jhonyy9:

from formula of area you can calcule the radius value and just substitute in place of radius inside formula of circumference hope helped

jhonyy9:

area =pi*r^2 400 = 3,14 *r^2 r^2 = 400/3,14 = 127,38 r = sqrt(127,38 = 11,28 cm yes you yre right with this

jhonyy9:

you are right sorry

jhonyy9:

and now just use this value of radius and calcule easy the circumference value

jhonyy9:

ok. ?

jhonyy9:

do you know formula for circumference ?

jhonyy9:

C = 2*pi*r

jhonyy9:

i need to go now it"s ok ?

jhonyy9:

lock ???

jhonyy9:

C=2*3,14*11,28 = ?

jhonyy9:

bye bye

lock:

thank you ! @jhonyy9